A couple is planning to finance its three-year-old sons university education. Money can be deposited at 5% compounded quartePlease do it asap
Answer #1

Required balance in savings a/c P+Px[1-(1:(1+r)^(n-1)]=r Here, 5.09% =(1+5%/4)^4-1 4 1 Interest rate per annum 2 Number of ye

Deposit in each quarter (P) FVA:([(1+r)^n-1]=r) Here, A Interest rate per annum B Number of years C Number of payments per pe

*Please rate thumbs up